### v1

@@ Abstract
This paper studies the linear quadratic regulation (LQR) problem of unknown discrete-time systems via dynamic output feedback learning control. In contrast to the state feedback, the optimality of the dynamic output feedback control for solving the LQR problem requires an implicit condition on the convergence of the state observer. Moreover, due to unknown system matrices and the existence of observer error, it is difficult to analyze the convergence and stability of most existing output feedback learning-based control methods. To tackle these issues, we propose a generalized dynamic output feedback learning control approach with guaranteed convergence, stability, and optimality performance for solving the LQR problem of unknown discrete-time linear systems. In particular, a dynamic output feedback controller is designed to be equivalent to a state feedback controller. This equivalence relationship is an inherent property without requiring convergence of the estimated state by the state observer, which plays a key role in establishing the off-policy learning control approaches. By value iteration and policy iteration schemes, the adaptive dynamic programming based learning control approaches are developed to estimate the optimal feedback control gain. In addition, a model-free stability criterion is provided by finding a nonsingular parameterization matrix, which contributes to establishing a switched iteration scheme. Furthermore, the convergence, stability, and optimality analyses of the proposed output feedback learning control approaches are given. Finally, the theoretical results are validated by two numerical examples.
@@ 1 Introduction
Reinforcement learning (RL) (Sutton & Barto, 2018; Lewis et al., 2012a), also known as adaptive dynamic programming (ADP) (Jiang & Jiang, 2017; Powell, 2004; Liu et al., 2021; Lewis et al., 2012b) in the field of control, is one of the most important cutting-edge research fields in artificial intelligence for optimization Wu et al. (2025) and decision-making in uncertain environments. Especially, machine learning method has shown promise in monitoring real-time green carbon dioxide tracking in refinery processes Cao et al. (2025). Moreover, the ADP algorithm offers iteration learning frameworks to estimate the optimal control policy for solving several optimal control problems by only accessing a series of historical data of input, output, and/or state instead of prior knowledge of system dynamics; see, Vrabie et al. (2009); Jiang & Jiang (2012); Modares et al. (2016); Chen et al. (2023); Modares & Lewis (2014); Gao & Jiang (2016); Yang et al. (2018); Rizvi & Lin (2020); Duan et al. (2024), and references therein. As an ongoing research topic, it is of great significance to advance the development of the ADP method with output feedback control, as it has the potential to effectively address the optimal control problem under the unmeasurable state in various practical settings. However, it is shown in Duan et al. (2024); Rizvi & Lin (2023); Chen et al. (2023) that the theoretical guarantees of the ADP-based output feedback learning control methods, including convergence, stability, and optimality, remain substantial challenges.
@@ 2.2 Plant and Preliminaries
Different from the state reconstruction given in (2.2) which is specific to the discrete-time systems, an alternative way to obtain the state information is to use a state observer to estimate the real state. Concretely, the state of system (1) can be estimated by using the following Luenburger observer:
@@
It is known that the Luenburger observer (10) has the property that $\lim_{k\rightarrow\infty}\hat{x}(k)=x(k)$ when all the eigenvalues of $(A-LC)$ are set within the unit circle under the observability condition. As given in Rizvi & Lin (2019), the estimated state $\hat{x}(k)$ can be parametrized by
@@
where the parameterization matrices $W_{y}\in\mathbb{R}^{n\times np}$ and $W_{u}\in\mathbb{R}^{n\times nm}$ are determined by the coefficients of the numerators in the transfer function matrix of the Luenburger observer system (10) with $y(k)$ and $u(k)$ being the inputs to the observer system, and the dynamics of $\omega(k)$ and $\sigma(k)$ are given as
@@ 3.1 Dynamic Output Feedback Controller Design
Proof. Consider a Luenburger observer system described as in (10) and rewiritten as
@@
$\varepsilon_{x}(k)$ $=\mathcal{Z}^{-1}\left(\frac{\varepsilon_{x}(0)}{(zI_{n}-A_{L})}\right)=\mathcal{Z}^{-1}\left(\frac{\mbox{adj}(zI_{n}-A_{L})\varepsilon_{x}(0)}{\mbox{det}(zI_{n}-A_{L})}\right)$
@@
$\hat{x}(k)$ $=\mathcal{Z}^{-1}\left(\sum_{i=1}^{m}\frac{B_{i}U_{i}(z)}{(zI_{n}-A_{L})}+\sum_{j=1}^{p}\frac{L_{j}Y_{j}(z)}{(zI_{n}-A_{L})}\right)$
@@ 3.2.1 Model-Free VI Scheme
Initialization. Choose a symmetric positive semi-definite matrix $\mathcal{P}_{0}$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain ${\mathcal{K}_{0}}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
@@ 3.2.2 Model-Free PI Scheme
Initialization. Choose an initial admissible control gain $\mathcal{K}_{0}$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
@@ 4.1 Convergence analysis
$=$ $\left[A_{L}^{n-1}\varepsilon_{0}+b_{1}A_{L}^{n-2}\varepsilon_{0}+\cdots+b_{n-1}\varepsilon_{0},~{}\dots,\right.$
@@ 4.2 Stability Analysis
Initialization. Define a utility function as $\mathcal{C}(k)$ in (2) and a modified utility function as $\tilde{\mathcal{C}}(k)$ in (69) with $Q_{\varepsilon}=Q_{\varepsilon}^{\mbox{\tiny T}}>0$. Select an pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain $\mathcal{K}_{0}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $\textbf{sign}_{\rm{PI}}=0$ and $j=0$.
@@ 4.3 Optimality Analysis
To analyze the optimality of several output feedback controllers with respect to the pre-defined value function (2), the optimality analysis among the proposed state parameterization based dynamic output feedback controller (37), the state reconstruction based dynamic output feedback controller (9), the Luenburger observer state parameterization based dynamic output feedback controller (12), and the static output feedback controller is given in this subsection.
@@
where $K^{*}$ is obtained by (5). The controller (12) equals to a traditional Luenburger state observer based output feedback controller
@@
where $\hat{x}(k)$ is the estimated state by the Luenburger state observer (10). This implies that the optimality analysis among controllers (9), (12), and (37) can be given by the analysis between the full-state feedback controller (74) and the Luenburger state observer based output feedback controller (75).
@@ 5.1 Unstable and uncontrollable system
The parameters in Algorithm 3 are chosen as $\mathcal{P}_{0}=I_{n_{z}}$ with $n_{z}=6$, $\mathcal{K}_{0}=\mathbf{0}$, and $\epsilon=0.001$. The exploration noise is a combination of $100$ sinusoids with random frequency $w_{i}$, amplitude $c_{i}$, and initial phase $d_{i}$, i.e., $\xi=\sum_{i=1}^{100}c_{i}\mbox{sin}(w_{i}k+d_{i})$. Two numerical simulations are carried out by employing Algorithm 3 with the proposed dynamic output feedback controller (denoted as $u$) and the traditional Luenburger observer-based output feedback controller (denoted as $\hat{u}$). The simulation results are shown in Figs. 33. Fig. 3 depicts the convergence performance of Algorithm 3, which shows that, after $83$ iterations, the estimated control gain is updated as
@@ 5.2 Application to the aircraft system
$\mathcal{A}_{\varepsilon}\!=\!\mathcal{A}\!=\!\left[\!\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -0.7786&-2.5391&-2.7600\end{array}\!\right],~{}~{}b\!=\!\left[\!\begin{array}[]{ccc}0\\ 0\\ 1\end{array}\!\right]$
@@
In addition, similar to Subsection 5.1, two numerical simulations are carried out by employing Algorithm 3 with $u$ and $\hat{u}$. The evolution of inputs and outputs of the closed-loop system is drawn in Fig. 9, which shows that the controllers are updated at $k_{N}=58$ when the rank condition is met. The trajectory of states under the proposed dynamic output feedback controller $u$ and the traditional Luenburger observer-based output feedback controller $\hat{u}$ are shown in Fig 9. The evolution of observer error during time instant $k\in[150,400]$ is shown in Fig. 9. From Figs. 99, we can see that the proposed output feedback learning control approach is immune to the observer error, while the transient performance of the closed-loop systems under the Luenburger observer-based output feedback controller is influenced by the observer error.

### v3

@@ Abstract
This paper studies the linear quadratic regulation (LQR) problem of unknown discrete-time systems via dynamic output feedback learning control. In contrast to the state feedback, the optimality of the dynamic output feedback control for solving the LQR problem requires an implicit condition on the convergence of the state observer. Moreover, due to unknown system matrices and the existence of observer error, it is difficult to analyze the convergence and stability of most existing output feedback learning-based control methods. To tackle these issues, we propose a generalized dynamic output feedback learning control approach with guaranteed convergence, stability, and optimality performance for solving the LQR problem of unknown discrete-time linear systems. In particular, a dynamic output feedback controller is designed to be equivalent to a state feedback controller. This equivalence relationship is an inherent property without requiring convergence of the estimated state by the state observer, which plays a key role in establishing the off-policy learning control approaches. By value iteration and policy iteration schemes, the adaptive dynamic programming based learning control approaches are developed to estimate the optimal feedback control gain. In addition, a model-free stability criterion is provided by finding a nonsingular parameterization matrix, which contributes to establishing a switched iteration scheme. Furthermore, the convergence, stability, and optimality analyses of the proposed output feedback learning control approaches are given. Finally, two examples are used to verify the effectiveness of the proposed dynamic output feedback learning control approach.
@@ 1 Introduction
Reinforcement learning (RL) (Sutton & Barto, 2018; Lewis et al., 2012a), also known as adaptive dynamic programming (ADP) (Jiang & Jiang, 2017; Powell, 2004; Liu et al., 2021; Lewis et al., 2012b) in the field of control, is one of the most important cutting-edge research fields in artificial intelligence for optimization Wu et al. (2025) and decision-making in uncertain environments. Especially, Reinforcement learning methods have shown promise in monitoring real-time green carbon dioxide tracking in refinery processes Cao et al. (2025) and in active traffic management Su et al. (2020). Moreover, the ADP algorithm offers iteration learning frameworks to estimate the optimal control policy for solving several optimal control problems by only accessing a series of historical data of input, output, and/or state instead of prior knowledge of system dynamics; see, Vrabie et al. (2009); Jiang & Jiang (2012); Modares et al. (2016); Chen et al. (2023); Modares & Lewis (2014); Gao & Jiang (2016); Yang et al. (2018); Rizvi & Lin (2020); Duan et al. (2024), and references therein. As an ongoing research topic, it is of great significance to advance the development of the ADP method with output feedback control, as it has the potential to effectively address the optimal control problem under the unmeasurable state in various practical settings. However, it is shown in Duan et al. (2024); Rizvi & Lin (2023); Chen et al. (2023) that the theoretical guarantees of the ADP-based output feedback learning control methods, including convergence, stability, and optimality, remain substantial challenges.
@@ 2.2 Plant and Preliminaries
Different from the state reconstruction given in (2.2) which is specific to the discrete-time systems, an alternative way to obtain the state information is to use a state observer to estimate the real state. Concretely, the state of system (1) can be estimated by using the following Luenberger observer:
@@
It is known that the Luenberger observer (10) has the property that $\lim_{k\rightarrow\infty}\hat{x}(k)=x(k)$ when all the eigenvalues of $(A-LC)$ are set within the unit circle under the observability condition. As given in Rizvi & Lin (2019), by setting $\hat{x}(k)=\bm{0}$, the estimated state $\hat{x}(k)$ can be parametrized by
@@
where the parameterization matrices $W_{y}\in\mathbb{R}^{n\times np}$ and $W_{u}\in\mathbb{R}^{n\times nm}$ are determined by the coefficients of the numerators in the transfer function matrix of the Luenberger observer system (10) with $y(k)$ and $u(k)$ being the inputs to the observer system, and the dynamics of $\omega(k)$ and $\sigma(k)$ are given as
@@ 3.1 Dynamic Output Feedback Controller Design
Proof. Consider a Luenberger observer system described as in (10) and rewiritten as
@@
$\varepsilon_{x}(k)$ $=\mathcal{Z}^{-1}\left((zI_{n}-A_{L})^{-1}{\varepsilon_{x}(0)}\right)$
$=\mathcal{Z}^{-1}\left(\frac{\mbox{adj}(zI_{n}-A_{L})\varepsilon_{x}(0)}{\mbox{det}(zI_{n}-A_{L})}\right)$
@@
$\hat{x}(k)$ $=\mathcal{Z}^{-1}\left(\sum_{i=1}^{m}{(zI_{n}-A_{L})}^{-1}{B_{i}U_{i}(z)}+\sum_{j=1}^{p}{(zI_{n}-A_{L})}^{-1}{L_{j}Y_{j}(z)}\right)$
@@ 3.2.1 Model-Free VI Scheme
Initialization. Choose a symmetric positive semi-definite matrix $\mathcal{P}_{0}$. Select a pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain ${\mathcal{K}_{0}}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
@@ 3.2.2 Model-Free PI Scheme
Initialization. Choose an initial admissible control gain $\mathcal{K}_{0}$. Select a pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $j=0$.
@@ 4.1 Convergence analysis
$=$ $\left[A_{L}^{n-1}\varepsilon_{0}+b_{1}A_{L}^{n-2}\varepsilon_{0}+\cdots+b_{n-1}\varepsilon_{0},\right.$
@@ 4.2 Stability Analysis
Initialization. Define a utility function as $\mathcal{C}(k)$ in (2) and a modified utility function as $\tilde{\mathcal{C}}(k)$ in (69) with $Q_{\varepsilon}=Q_{\varepsilon}^{\mbox{\tiny T}}>0$. Select a pair $(\mathcal{G}_{1},\mathcal{G}_{2})$, an initial internal state $\eta(0)$, an initial control gain $\mathcal{K}_{0}$, and a small threshold $\epsilon>0$. Apply the initial control policy $u_{0}=-{\mathcal{K}_{0}}{\eta}+\xi$ with exploration noise $\xi$. Set $\textbf{sign}_{\rm{PI}}=0$ and $j=0$.
@@ 4.3 Optimality Analysis
To analyze the optimality of several output feedback controllers with respect to the pre-defined value function (2), the optimality analysis among the proposed state parameterization based dynamic output feedback controller (37), the state reconstruction based dynamic output feedback controller (9), the Luenberger observer state parameterization based dynamic output feedback controller (12), and the static output feedback controller is given in this subsection.
@@
where $K^{*}$ is obtained by (5). The controller (12) equals to a traditional Luenberger state observer based output feedback controller
@@
where $\hat{x}(k)$ is the estimated state by the Luenberger state observer (10). This implies that the optimality analysis among controllers (9), (12), and (37) can be given by the analysis between the full-state feedback controller (74) and the Luenberger state observer based output feedback controller (75).
@@ 5.1 Unstable and uncontrollable system
The parameters in Algorithm 3 are chosen as $\mathcal{P}_{0}=I_{n_{z}}$ with $n_{z}=6$, $\mathcal{K}_{0}=\mathbf{0}$, and $\epsilon=0.001$. The exploration noise is a combination of $100$ sinusoids with random frequency $w_{i}$, amplitude $c_{i}$, and initial phase $d_{i}$, i.e., $\xi=\sum_{i=1}^{100}c_{i}\mbox{sin}(w_{i}k+d_{i})$. Two numerical simulations are carried out by employing Algorithm 3 with the proposed dynamic output feedback controller (denoted as $u$) and the traditional Luenberger observer-based output feedback controller (denoted as $\hat{u}$). The simulation results are shown in Figs. 33. Fig. 3 depicts the convergence performance of Algorithm 3, which shows that, after $83$ iterations, the estimated control gain is updated as
@@ 5.2 Application to the aircraft system
$\mathcal{A}_{\varepsilon}$ $\!=\left[\!\begin{array}[]{ccc}-0.91&1.1&-1.2\\ 0&-0.92&~{}~{}~{}\!1.3\\ 0&0&-0.93\end{array}\!\right]$
$\mathcal{A}$ $\!=\left[\!\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -0.7786&-2.5391&-2.7600\end{array}\!\right],~{}~{}b\!=\!\left[\!\begin{array}[]{ccc}0\\ 0\\ 1\end{array}\!\right]$
@@
In addition, similar to Subsection 5.1, two numerical simulations are carried out by employing Algorithm 3 with $u$ and $\hat{u}$. The evolution of inputs and outputs of the closed-loop system is drawn in Fig. 9, which shows that the controllers are updated at $k_{N}=58$ when the rank condition is met. The trajectory of states under the proposed dynamic output feedback controller $u$ and the traditional Luenberger observer-based output feedback controller $\hat{u}$ are shown in Fig 9. The evolution of observer error during time instant $k\in[150,400]$ is shown in Fig. 9. From Figs. 99, we can see that the proposed output feedback learning control approach is immune to the observer error, while the transient performance of the closed-loop systems under the Luenberger observer-based output feedback controller is influenced by the observer error.